# Strict comparison in reduced group $C^{*}$-algebras

### v1

@@ 1.2. Context and main results
A prototypical example of a reduced group $C^{*}$-algebra is the one arising from the countable free group $\mathbb{F}_{2}$. The study of this $C^{*}$-algebra has been extremely important for opening up several deep insights in the subject. Some examples of this include the following fundamental results: unique trace and simplicity [50] (see also the major works [37, 33] which were inspired by Powers’ insights), lacking non-trivial projections [48], $K$-theory computations [48], and showing stable rank 1 [20] in the sense of Rieffel [52]. Despite these important developments, the fundamental question of proving strict comparison for $C^{*}_{r}(\mathbb{F}_{2})$ has evaded experts. Within the community, it has been a widely-known open problem with initial motivation stemming from the work of Dykema-Rørdam in 1998 ([22, 23], see also [53, Proposition 6.3.2]), which addresses comparison in $C^{*}_{r}(\mathbb{F}_{\infty})$. Note that the ideas behind the proof for strict comparison for $C^{*}_{r}(\mathbb{F}_{\infty})$ cannot be readily used to prove this result for $C^{*}_{r}(\mathbb{F}_{2})$. For instance, the proof there uses the ability to find freely independent elements from any finite subsets, up to norm perturbations, which holds inside $C^{*}_{r}(\mathbb{F}_{\infty})$. However such behavior is not clear at all to witness inside $C^{*}_{r}(\mathbb{F}_{2})$. Strict comparison for $C^{*}_{r}({\mathbb{F}}_{2})$ has been explicitly stated as an open question in the work of Robert in 2012 [53] and in the recent substantial problem list in $C^{*}$-algebras [58], see Problems 86 and 88.
@@
Thanks to major developments [55, 43], we now know that for unital, separable, simple, unique trace, nuclear, infinite-dimensional $C^{*}$-algebras, strict comparison and $\mathcal{Z}$-stability are equivalent. Importantly, at the moment, outside the $\mathcal{Z}$-stable case there are only a few examples of strict comparison arising from groups, such as infinite reduced free products like $C^{*}_{r}(\mathbb{F}_{\infty})$ ([53, Proposition 6.3.2]). With the next main result, we greatly expand the list of known examples. To motivate this, we need to discuss the following notion. We say a group $G$ has the rapid decay property if, for every finite generating set $X$, there exists a polynomial $P$ such that for every element $a\in\mathbb{C}G$ supported on $B_{X}(r)$, $\left\|\lambda_{G}(a)\right\|\leq P(r)\|a\|_{2}.$ The rapid decay property (see the survey [56]) was first proved by Haagerup for the free groups [31]. It was then generalized to Gromov hyperbolic groups [36, 34], relatively hyperbolic groups with peripheral subgroups having rapid decay [56], cocompact lattices in $SL_{3}(\mathbb{R})$ or $SL_{3}(\mathbb{C})$ [40, 12], mapping class groups of surfaces [5], large type Artin groups [13], and more. Rapid decay is extremely influential and has been used to prove several important problems, including the Novikov conjecture [14], the Baum-Connes conjecture [41], and others including [30]. Before we state our result, we will need to recall one more important and influential notion from geometric group theory [47]. A group is said to be acylindrically hyperbolic if it admits a non-elementary acylindrical action on a Gromov hyperbolic metric space. We direct the reader to [46] for context, examples, and references. We are now ready to state our next result:
@@
Note that the trivial finite radical assumption in the above theorem is necessary due to the fact that these $C^{*}$-algebras will have unique trace as a consequence of being selfless (see also [9]). The following are some examples of groups (with an added trivial finite radical assumption) that fit into the above result (see [46, 18]): non-elementary hyperbolic groups; non-elementary relatively hyperbolic groups relative to peripheral subgroups with the rapid decay property, which includes free products of groups with the rapid decay property; mapping class groups of compact orientable surfaces [6]; graph products of rapid decay groups over graphs without a splitting [44]; etc. We speculate that our proof technique might open up a way to prove strict comparison for even wider families of groups with rapid decay, such as all $C^{*}$-simple groups, see Remark 3.6.
Importantly, see Remark 3.9, which describes why the rapid decay assumption cannot be removed with our current approach. However, we point out that our result gives new examples of groups which do not have repaid decay property but whose reduced $C^{*}$-algebras are selfless. Indeed by Theorem A and Theorem 4.2 of [54], we have $C^{*}_{r}(\mathbb{F}_{2}*SL_{3}(\mathbb{Z}))$ is selfless, however $\mathbb{F}_{2}*SL_{3}(\mathbb{Z})$ does not have rapid decay.
@@ 1.3. Applications to classification and Cuntz semigroup theory
It is a very standard fact that every II1 factor contains a copy of the hyperfinite II1 factor $R$ uniquely up to pointwise approximate unitary equivalence. The most naïve analogue of this in $C^{*}$-algebras would be that every simple, unital, infinite-dimensional $C^{*}$-algebra contains the Jiang-Su algebra $\mathcal{Z}$ uniquely up to pointwise approximate unitary equivalence. Both existence and uniqueness problems for embeddings of $\mathcal{Z}$ tend to be difficult, see [58, Section 18]. It is known that there are simple, unital, infinite-dimensional $C^{*}$-algebras that do not contain $\mathcal{Z}$ [17]. It remains open if there is any unital $C^{*}$-algebra $A$ with two non-equivalent unital embedding of $\mathcal{Z}$. Our work yields the following concrete application. The result below follows from Proposition 6.3.1 in [53] and our Theorem B. Note that for our examples, the uniqueness of the trace follows from the fact that the groups in consideration have trivial amenable radical and the unique quasitrace comes from exactness and Haagerup’s famous quasitrace theorem [32]. Furthermore, stable rank one holds for all acylindrically hyperbolic groups [26]. Note also that the following result addresses problems 62 and 64 in [58]:
@@
Note that, our Theorem B can be used to fully compute the Cuntz semigroup for the vast family of acylindrically hyperbolic groups with trivial finite radical and with rapid decap property as described above. This partially settles the general problem in the paragraph before Problem 16.4 in [25], obtaining a full computation for many classes of $C^{*}$-simple groups. For examples of groups with stable rank one, see [20, 26, 51]. Note also that Theorem B additionally recovers with a new proof stable rank one for the families of groups in the statement.
@@ 1.5. Comments
The proofs of these results split into a hands-on computational part (showing selflessness in our sense, see 3.1) and a soft $C^{*}$-algebra part (using rapid decay à la Louder-Magee, see Theorem 3.5). If the reader would like to just focus on understanding the proof of Theorem A, we direct them to a self contained approach via the combination of Proposition 3.2 and Theorem 3.5. The quantitative part of the proof of Theorem B is rather involved, but quite natural from a geometric point of view. Appendix A contains a quantitative version of a key theorem of Wenyuan Yang that is needed to prove our result B. This is quite technical and we recommend the interested reader to side by side refer to [66, Corollary 3.4].
@@ Acknowledgments
This result was proved during the visit of the first author to UCSD in Fall 2024, supported by the third author’s NSF grant DMS 2350049 and the Institute of Mathematics of the Polish Academy of Sciences. The second and fourth author were supported in part by NSF grant DMS 2153805. We thank our colleagues J. Gabe, D. Jekel, N.C. Phillips, L. Robert, M. Rørdam, C. Schafhauser, G. Szabo, H. Thiel and S. White for extremely helpful comments on the introduction. We sincerely thank all the experts we shared our earlier draft with, for providing us great encouragement and helpful comments to improve our paper. We also thank our colleagues F. Fournier-Facio, J. Huang, Y. Lodha, D. Osin, A. Sisto, and W. Yang for helpful pointers.
@@ 3.2. A general family of selfless groups
Fix a finite generating set $S$ for $G$. Since $G$ is acylindrically hyperbolic and has trivial finite radical, by [18, Theorem 6.14], there is a hyperbolically embedded copy of ${\mathbb{F}}_{2}$ in $G$. Denote by $a,b$ the generators of ${\mathbb{F}}_{2}$. Since $\langle a\rangle$ is hyperbolically embedded in ${\mathbb{F}}_{2}$, by transitivity of hyperbolic embeddings [18, Proposition 4.35] we have that $\langle a\rangle$ is hyperbolically embedded in $G$. Therefore $\langle a\rangle$ is almost malnormal in $G$ [18, Proposition 4.33], implying that the maximal virtually cyclic subgroup containing $a,$ namely $E(a)$, is equal to $\langle a\rangle$ (see [24] for details on virtually cyclic groups). By [18, Theorem 4.31], we have that this copy of ${\mathbb{F}}_{2}$, with word metric given by its generating set $\{a,b\}$, is a Lipschitz quasi-retract of $G$ with word metric given by $S$. Therefore there are constants $C,D$ such that whenever $g\in{\mathbb{F}}_{2},$ $|g|_{\{a,b\}}\leq C|g|_{S}$ and $|g|_{S}\leq D|g|_{\{a,b\}}$. Since $a$ has infinite order, [46, Theorem 1.4] says that there is an acylindrical action of $G$ on a hyperbolic space $X$ such that $a$ is loxodromic.
@@ 3.3. Proving strict comparison for $C^{*}_{r}(G)$
We follow the proof of [42, Theorem 1.6]. Set $Y=X\cup\{a\}$ to be our finite generating set of $\Gamma:=G*{\mathbb{Z}}$. Let $\varepsilon>0$ and let $S\subset\Gamma$ be a finite subset. Then there is some $R>0$ such that $S\subset B_{Y}(R)$. Let $S_{Y}(R)\subset{\mathbb{C}}\Gamma$ denote the elements supported on $B_{Y}(R)$ with $\ell^{1}$ norm equal to 1. Using compactness of $S_{Y}(R)$ with respect to the $\ell^{1}$ norm, we find a finite $\frac{\varepsilon}{3}$-net $\{a_{i}\}_{i\in I}$ for $S_{Y}(R)$. It now suffices to prove the existence of a homomorphism $\phi:\Gamma\to G$ such that $\phi|_{G}=\operatorname{id}_{G}$ and that, for all $i\in I,$
@@
By [54, Proposition 5], it suffices to show that, for some free ultrafilter ${\mathcal{U}}$ on ${\mathbb{N}}$ there is an injective $\ast$-homomorphism $\phi$ from the reduced free product $C^{*}_{r}(G)*C_{r}^{*}({\mathbb{Z}})$ to $C^{*}_{r}(G)^{\mathcal{U}}$ such that, composing with the canonical inclusion $C_{r}^{*}(G)\subset C_{r}^{*}(G)*C_{r}^{*}({\mathbb{Z}})$, we obtain the diagonal embedding $C_{r}^{*}(G)\subset C_{r}^{*}(G)^{\mathcal{U}}$. In other words, we require an existential embedding of $C^{*}_{r}(G)*C_{r}^{*}({\mathbb{Z}})$ in $C^{*}_{r}(G)^{\mathcal{U}}$. But this is precisely the content of Proposition 3.7. ∎
@@ Appendix A Quantitative version of Wenyuan Yang’s admissible path lemma
The aim of this section is to prove a quantitative version of W. Yang’s admissible path lemma [66, Corollary 3.4], which applies to our setting. For our purposes, we will take a fixed action on a hyperbolic space but vary the loxodromic element and thus the quasi-geodesic constants $\lambda$. So throughout, we track dependence on $\lambda$. The following quantitative version of the Morse lemma is present in [27, Section 5.2]; see [60, Theorem 1.1]) for an explicit polynomial.
@@
([19, Proposition 5]) Let $G\curvearrowright X$ be an acylindrical action on a hyperbolic space and let $g\in G$ be a loxodromic element (i.e. a hyperbolic isometry). Then for all $h\not\in E(g)$,
@@
The following lemma is the quantitative version of [62, Lemma 2.23].
@@
The following theorem, the main theorem of this section, is essentially a reproduction of [66, Corollary 3.4] but where we have tracked all of the constants and dependencies to verify they grow polynomially.
@@
Set $\lambda=\frac{d(go,o)}{\tau(g)}$. Let ${\mathcal{L}}$ be the collection of $(\lambda,0)$-quasi-geodesics in $X$; in particular, ${\mathcal{L}}$ includes all geodesics since (without loss of generality) $\lambda\geq 1$. Let ${\mathbb{X}}$ be defined by ${\mathbb{X}}=\{\cup_{n=k}^{\ell}[hg^{n},0,hg^{n+1}o]:h\not\in E(g)\}$ and note that ${\mathbb{X}}\subset{\mathcal{L}}.$ By Lemma A.1 and [7, Section 3], we have that ${\mathbb{X}}$ is $(\mu,\varepsilon)$-contracting with respect to ${\mathcal{L}}$ where $\mu,\varepsilon$ are uniformly bounded above by a quadratic $Q_{3}(\lambda)$. That is, whenever $q\in{\mathcal{L}}$ and $p\in{\mathbb{X}}$ such that $d(q,p)\geq Q_{3}(\lambda)$ then $\|\pi_{p}(q)\|<Q_{3}(\lambda)$. (In the notation of [66], $\mu(\lambda^{\prime},c^{\prime}),\varepsilon(\lambda^{\prime},c^{\prime})\leq Q_{3}(\lambda)$ for all $\lambda^{\prime},c^{\prime}\in{\mathbb{R}}$.)
@@
By Lemma A.3, if $p,p^{\prime}\in{\mathbb{X}}$ are quasi-geodesics corresponding to $h,h^{\prime}$ in different left cosets of $E(g)$, then $p,p^{\prime}$ have bounded projection $T$ where $T=Q_{2}(\lambda)(1+[g])=O(\lambda^{3})$ since $[g]\leq d(go,o)=\tau(g)\lambda$ (and the stable length $\tau(g)$ is fixed). We note that in [66], the symbol $\tau$ is used for bounded projection but we have already used $\tau$ for stable length. By [66, Lemma 2.7], $p,p^{\prime}$ also have $\nu$-bounded intersection where $\nu(U)=2U+O(\lambda^{3})$.
@@
We are now ready to analyze the constants in the proof of [66, Proposition 3.3]. We have $c=0$ everywhere.
@@
Now, if we choose all of the $n_{i}$ so that $|n_{i}|$ are larger than $\lambda D=O(|g|_{S}^{6}),$ we can apply [66, Corollary 3.4] to obtain that our path $P$ is a $(\Lambda,0)$-quasi-geodesic. In particular, $P$ is nontrivial and so it must be that $h_{1}g^{n_{1}}h_{2}g^{n_{2}}\cdots h_{m}g^{n_{m}}\neq e$. (Note that in the case $n_{m}=0$, we can still apply [66, Corollary 3.4].) This finishes the proof. ∎

### v2

@@ 1.2. Context and main results
A prototypical example of a reduced group $C^{*}$-algebra is the one arising from the countable free group $\mathbb{F}_{2}$. The study of this $C^{*}$-algebra has been extremely important for opening up several deep insights in the subject. Some examples of this include the following fundamental results: unique trace and simplicity [52] (see also the major works [39, 35] which were inspired by Powers’ insights), lacking non-trivial projections [50], $K$-theory computations [50], and showing stable rank 1 [22] in the sense of Rieffel [54]. Despite these important developments, the fundamental question of proving strict comparison for $C^{*}_{r}(\mathbb{F}_{2})$ has evaded experts. Within the community, it has been a widely-known open problem with initial motivation stemming from the work of Dykema-Rørdam in 1998 ([24, 25], see also [55, Proposition 6.3.2]), which addresses comparison in $C^{*}_{r}(\mathbb{F}_{\infty})$. Note that the ideas behind the proof for strict comparison for $C^{*}_{r}(\mathbb{F}_{\infty})$ cannot be readily used to prove this result for $C^{*}_{r}(\mathbb{F}_{2})$. For instance, the proof there uses the ability to find freely independent elements from any finite subsets, up to norm perturbations, which holds inside $C^{*}_{r}(\mathbb{F}_{\infty})$. However such behavior is not clear at all to witness inside $C^{*}_{r}(\mathbb{F}_{2})$. Strict comparison for $C^{*}_{r}({\mathbb{F}}_{2})$ has been explicitly stated as an open question in the work of Robert in 2012 [55] and in the recent substantial problem list in $C^{*}$-algebras [60], see Problems 86 and 88.
@@
Thanks to major developments [57, 45], we now know that for unital, separable, simple, unique trace, nuclear, infinite-dimensional $C^{*}$-algebras, strict comparison and $\mathcal{Z}$-stability are equivalent. Importantly, at the moment, outside the $\mathcal{Z}$-stable case there are only a few examples of strict comparison arising from groups, such as infinite reduced free products like $C^{*}_{r}(\mathbb{F}_{\infty})$ ([55, Proposition 6.3.2]). With the next main result, we greatly expand the list of known examples. To motivate this, we need to discuss the following notion. We say a group $G$ has the rapid decay property if, for every finite generating set $X$, there exists a polynomial $P$ such that for every element $a\in\mathbb{C}G$ supported on $B_{X}(r)$, $\left\|\lambda_{G}(a)\right\|\leq P(r)\|a\|_{2}.$ The rapid decay property (see the survey [58]) was first proved by Haagerup for the free groups [33]. It was then generalized to Gromov hyperbolic groups [38, 36], relatively hyperbolic groups with peripheral subgroups having rapid decay [58], cocompact lattices in $SL_{3}(\mathbb{R})$ or $SL_{3}(\mathbb{C})$ [42, 14], mapping class groups of surfaces [7], large type Artin groups [15], heirarchically hyperbolic groups [5], and more. Rapid decay is extremely influential and has been used to prove several important problems, including the Novikov conjecture [16], the Baum-Connes conjecture [43], and others including [32]. Before we state our result, we will need to recall one more important and influential notion from geometric group theory [49]. A group is said to be acylindrically hyperbolic if it admits a non-elementary acylindrical action on a Gromov hyperbolic metric space. We direct the reader to [48] for context, examples, and references. We are now ready to state our next result:
@@
Note that the trivial finite radical assumption in the above theorem is necessary due to the fact that these $C^{*}$-algebras will have unique trace as a consequence of being selfless (see also [11]). The following are some examples of groups (with an added trivial finite radical assumption) that fit into the above result (see [48, 20]): non-elementary hyperbolic groups; non-elementary relatively hyperbolic groups relative to peripheral subgroups with the rapid decay property, which includes free products of groups with the rapid decay property; mapping class groups of compact orientable surfaces [8]; graph products of rapid decay groups over graphs without a splitting [46]; heirarchically hyperbolic groups [6]. In light of [42], it is very timely and interesting to ask if our main results hold in the setting of higher rank lattices, especially cocompact lattices in $SL_{3}(\mathbb{R})$. Our proof strategy has no bearing in this setting, and therefore we leave it as an important open question arising from our work. Notably, even stable rank $1$ seems to be an open problem for reduced group $C^{*}$-algebras of higher rank lattices.
Importantly, also see Remark 3.9, which describes why the rapid decay assumption cannot be removed with our current approach. However, we point out that our result gives new examples of groups which do not have rapid decay property but whose reduced $C^{*}$-algebras are selfless. Indeed by Theorem A and Theorem 4.2 of [56], we have $C^{*}_{r}(\mathbb{F}_{2}*SL_{3}(\mathbb{Z}))$ is selfless, however $\mathbb{F}_{2}*SL_{3}(\mathbb{Z})$ does not have rapid decay.
@@ 1.3. Applications to classification and Cuntz semigroup theory
It is a very standard fact that every II1 factor contains a copy of the hyperfinite II1 factor $R$ uniquely up to pointwise approximate unitary equivalence. The most naïve analogue of this in $C^{*}$-algebras would be that every simple, unital, infinite-dimensional $C^{*}$-algebra contains the Jiang-Su algebra $\mathcal{Z}$ uniquely up to pointwise approximate unitary equivalence. Both existence and uniqueness problems for embeddings of $\mathcal{Z}$ tend to be difficult, see [60, Section 18]. It is known that there are simple, unital, infinite-dimensional $C^{*}$-algebras that do not contain $\mathcal{Z}$ [19]. It remains open if there is any unital $C^{*}$-algebra $A$ with two non-equivalent unital embedding of $\mathcal{Z}$. Our work yields the following concrete application. The result below follows from Proposition 6.3.1 in [55] and our Theorem B. Note that for our examples, the uniqueness of the trace follows from the fact that the groups in consideration have trivial amenable radical and the unique quasitrace comes from exactness and Haagerup’s famous quasitrace theorem [34]. Furthermore, stable rank one holds for all acylindrically hyperbolic groups [28]. Note also that the following result addresses problems 62 and 64 in [60]:
@@
Note that, our Theorem B can be used to fully compute the Cuntz semigroup for the vast family of acylindrically hyperbolic groups with trivial finite radical and with rapid decap property as described above. This partially settles the general problem in the paragraph before Problem 16.4 in [27], obtaining a full computation for many classes of $C^{*}$-simple groups. For examples of groups with stable rank one, see [22, 28, 53]. Note also that Theorem B additionally recovers with a new proof stable rank one for the families of groups in the statement (via [56]).
@@ 1.5. Comments
The proofs of these results split into a hands-on computational part (showing selflessness in our sense, see 3.1) and a soft $C^{*}$-algebra part (using rapid decay à la Louder-Magee, see Theorem 3.5). If the reader would like to just focus on understanding the proof of Theorem A, we direct them to a self contained approach via the combination of Proposition 3.2 and Theorem 3.5. The quantitative part of the proof of Theorem B is rather involved, but quite natural from a geometric point of view. Appendix A contains a quantitative version of a key theorem of Wenyuan Wang that is needed to prove our result B. This is quite technical and we recommend the interested reader to side by side refer to [68, Corollary 3.4].
@@ Acknowledgments
This result was proved during the visit of the first author to UCSD in Fall 2024, supported by the third author’s NSF grant DMS 2350049 and the Institute of Mathematics of the Polish Academy of Sciences. The second and fourth author were supported in part by NSF grant DMS 2153805. We thank our colleagues J. Gabe, D. Jekel, N.C. Phillips, L. Robert, M. Rørdam, C. Schafhauser, G. Szabo, H. Thiel and S. White for extremely helpful comments on the introduction. We sincerely thank all the experts we shared our earlier draft with, for providing us great encouragement and helpful comments to improve our paper. We also thank our colleagues J. Behrstock, F. Fournier-Facio, J. Huang, Y. Lodha, D. Osin, A. Sisto, and W. Yang for helpful pointers.
@@ 3.2. A general family of selfless groups
Fix a finite generating set $S$ for $G$. Since $G$ is acylindrically hyperbolic and has trivial finite radical, by [20, Theorem 6.14], there is a hyperbolically embedded copy of ${\mathbb{F}}_{2}$ in $G$. Denote by $a,b$ the generators of ${\mathbb{F}}_{2}$. Since $\langle a\rangle$ is hyperbolically embedded in ${\mathbb{F}}_{2}$, by transitivity of hyperbolic embeddings [20, Proposition 4.35] we have that $\langle a\rangle$ is hyperbolically embedded in $G$. Therefore $\langle a\rangle$ is almost malnormal in $G$ [20, Proposition 4.33], implying that the maximal virtually cyclic subgroup containing $a,$ namely $E(a)$, is equal to $\langle a\rangle$ (see [26] for details on virtually cyclic groups). By [20, Theorem 4.31], we have that this copy of ${\mathbb{F}}_{2}$, with word metric given by its generating set $\{a,b\}$, is a Lipschitz quasi-retract of $G$ with word metric given by $S$. Therefore there are constants $C,D$ such that whenever $g\in{\mathbb{F}}_{2},$ $|g|_{\{a,b\}}\leq C|g|_{S}$ and $|g|_{S}\leq D|g|_{\{a,b\}}$. Since $a$ has infinite order, [48, Theorem 1.4] says that there is an acylindrical action of $G$ on a hyperbolic space $X$ such that $a$ is loxodromic.
@@ 3.3. Proving strict comparison for $C^{*}_{r}(G)$
We follow the proof of [44, Theorem 1.6]. Set $Y=X\cup\{a\}$ to be our finite generating set of $\Gamma:=G*{\mathbb{Z}}$. Let $\varepsilon>0$ and let $S\subset\Gamma$ be a finite subset. Then there is some $R>0$ such that $S\subset B_{Y}(R)$. Let $S_{Y}(R)\subset{\mathbb{C}}\Gamma$ denote the elements supported on $B_{Y}(R)$ with $\ell^{1}$ norm equal to 1. Using compactness of $S_{Y}(R)$ with respect to the $\ell^{1}$ norm, we find a finite $\frac{\varepsilon}{3}$-net $\{a_{i}\}_{i\in I}$ for $S_{Y}(R)$. It now suffices to prove the existence of a homomorphism $\phi:\Gamma\to G$ such that $\phi|_{G}=\operatorname{id}_{G}$ and that, for all $i\in I,$
@@
By [56, Proposition 5], it suffices to show that, for some free ultrafilter ${\mathcal{U}}$ on ${\mathbb{N}}$ there is an injective $\ast$-homomorphism $\phi$ from the reduced free product $C^{*}_{r}(G)*C_{r}^{*}({\mathbb{Z}})$ to $C^{*}_{r}(G)^{\mathcal{U}}$ such that, composing with the canonical inclusion $C_{r}^{*}(G)\subset C_{r}^{*}(G)*C_{r}^{*}({\mathbb{Z}})$, we obtain the diagonal embedding $C_{r}^{*}(G)\subset C_{r}^{*}(G)^{\mathcal{U}}$. In other words, we require an existential embedding of $C^{*}_{r}(G)*C_{r}^{*}({\mathbb{Z}})$ in $C^{*}_{r}(G)^{\mathcal{U}}$. But this is precisely the content of Proposition 3.7. ∎
@@ Appendix A Quantitative version of Wenyuan Yang’s admissible path lemma
The aim of this section is to prove a quantitative version of W. Yang’s admissible path lemma [68, Corollary 3.4], which applies to our setting. For our purposes, we will take a fixed action on a hyperbolic space but vary the loxodromic element and thus the quasi-geodesic constants $\lambda$. So throughout, we track dependence on $\lambda$. The following quantitative version of the Morse lemma is present in [29, Section 5.2]; see [62, Theorem 1.1]) for an explicit polynomial.
@@
([21, Proposition 5]) Let $G\curvearrowright X$ be an acylindrical action on a hyperbolic space and let $g\in G$ be a loxodromic element (i.e. a hyperbolic isometry). Then for all $h\not\in E(g)$,
@@
The following lemma is the quantitative version of [64, Lemma 2.23].
@@
The following theorem, the main theorem of this section, is essentially a reproduction of [68, Corollary 3.4] but where we have tracked all of the constants and dependencies to verify they grow polynomially.
@@
Set $\lambda=\frac{d(go,o)}{\tau(g)}$. Let ${\mathcal{L}}$ be the collection of $(\lambda,0)$-quasi-geodesics in $X$; in particular, ${\mathcal{L}}$ includes all geodesics since (without loss of generality) $\lambda\geq 1$. Let ${\mathbb{X}}$ be defined by ${\mathbb{X}}=\{\cup_{n=k}^{\ell}[hg^{n},0,hg^{n+1}o]:h\not\in E(g)\}$ and note that ${\mathbb{X}}\subset{\mathcal{L}}.$ By Lemma A.1 and [9, Section 3], we have that ${\mathbb{X}}$ is $(\mu,\varepsilon)$-contracting with respect to ${\mathcal{L}}$ where $\mu,\varepsilon$ are uniformly bounded above by a quadratic $Q_{3}(\lambda)$. That is, whenever $q\in{\mathcal{L}}$ and $p\in{\mathbb{X}}$ such that $d(q,p)\geq Q_{3}(\lambda)$ then $\|\pi_{p}(q)\|<Q_{3}(\lambda)$. (In the notation of [68], $\mu(\lambda^{\prime},c^{\prime}),\varepsilon(\lambda^{\prime},c^{\prime})\leq Q_{3}(\lambda)$ for all $\lambda^{\prime},c^{\prime}\in{\mathbb{R}}$.)
@@
By Lemma A.3, if $p,p^{\prime}\in{\mathbb{X}}$ are quasi-geodesics corresponding to $h,h^{\prime}$ in different left cosets of $E(g)$, then $p,p^{\prime}$ have bounded projection $T$ where $T=Q_{2}(\lambda)(1+[g])=O(\lambda^{3})$ since $[g]\leq d(go,o)=\tau(g)\lambda$ (and the stable length $\tau(g)$ is fixed). We note that in [68], the symbol $\tau$ is used for bounded projection but we have already used $\tau$ for stable length. By [68, Lemma 2.7], $p,p^{\prime}$ also have $\nu$-bounded intersection where $\nu(U)=2U+O(\lambda^{3})$.
@@
We are now ready to analyze the constants in the proof of [68, Proposition 3.3]. We have $c=0$ everywhere.
@@
Now, if we choose all of the $n_{i}$ so that $|n_{i}|$ are larger than $\lambda D=O(|g|_{S}^{6}),$ we can apply [68, Corollary 3.4] to obtain that our path $P$ is a $(\Lambda,0)$-quasi-geodesic. In particular, $P$ is nontrivial and so it must be that $h_{1}g^{n_{1}}h_{2}g^{n_{2}}\cdots h_{m}g^{n_{m}}\neq e$. (Note that in the case $n_{m}=0$, we can still apply [68, Corollary 3.4].) This finishes the proof. ∎